OpenStudy (anonymous):

Find the exact value of tan[cos^-1(1/3)]

OpenStudy (jdoe0001):

\(\bf cos^{-1}\left(\frac{1}{3}\right)=\theta\implies cos(\theta)=\cfrac{1}{3}\to \cfrac{adjacent}{hypotenuse}\to \cfrac{a=1}{c=3}\)  \(\bf recall\implies tan(\theta)=\cfrac{opposite}{hypotenuse}\to \cfrac{b}{c}\)